# Eliciting Latent Predictions from Transformers with the Tuned Lens

### v4

@@ 1 Introduction
We use the tuned lens to gain qualitative insight into the computational process of pretrained language models, by examining how their latent predictions evolve during a forward pass (Figure 1, Appendix B)
@@ 2 The Logit Lens
$\mathcal{M}_{>\ell}(\boldsymbol{h}_{\ell})=\mathrm{LayerNorm}\Big{[}\boldsymbol{h}_{\ell}+\sum_{\ell^{\prime}=\ell}^{L}\underbrace{F_{\ell^{\prime}}(\boldsymbol{h}_{\ell^{\prime}})}_{\text{residual update}}\Big{]}W_{U}.$ (2)
@@
Moreover, the type of information extracted by the logit lens varies both from model to model and from layer to layer, making it difficult to interpret. For example, we find that for BLOOM and OPT 125M, the top 1 prediction of the logit lens is often the input token, rather than any plausible continuation token, in more than half the layers (Figure 18).
@@
[…]
@@ 3 The Tuned Lens
Figure 6: Pairwise similarities of hidden state covariance matrices across layers of Pythia 12B. Layer 4 introduces two outlier dimensions which dominate the covariance; removing them reveals smooth representational drift with depth. To control for varying hidden state norms, we measure the Frobenius cosine similarity, or $\frac{\langle A,B\rangle_{F}}{\|A\|_{F}\|B\|_{F}}$ for two matrices $A$ and $B$.
Results. We plot tuned lens perplexity as a function of depth for the Pythia models and GPT-NeoX-20B in Figure 5333Pythia and GPT-NeoX-20B were trained using the same architecture, data, and codebase (Andonian et al., 2021). While they’re not officially the same model suite, they’re more consistent than the OPT models.; results for other model families can be found in Appendix A.
We find that the tuned lens resolves the problems with the logit lens discussed in Section 2: it has significantly lower bias (Figure 3), and much lower perplexity than the logit lens across the board (Figure 5, Appendix A).
@@
We report transfer penalties for the largest Pythia model in Figure 7. Overall, transfer penalties are quite low, especially for nearby layers (entries near the diagonal in Figure 7). Comparing to the two plots in Figure 6, we notice that transfer penalties are strongly negatively correlated with covariance similarity (Spearman $\rho=-0.78$). Unlike Figure 6, however, Figure 7 is not symmetric: transfer penalties are higher when training on a layer with the outlier dimensions (Layer 5 and later) and testing on a layer without them, than the reverse.
@@
As an example of the transferability of lenses to fine-tuned models we use Vicuna 13B (Chiang et al., 2023), an open source instruct fine-tuned chat model based on LLaMA 13B (Touvron et al., 2023). In Figure 12, Appendix A, we compare the performance of a lens specifically trained on Vicuna to a tuned lens trained only on LLaMA. Both Tuned Lenses were trained using a subsample of the RedPajama dataset (Together, 2023), an open source replication of LLaMA’s training set. The performance of both lenses was then evaluated on the test set of Anthropic’s Helpful Harmless (Bai et al., 2022) conversation dataset and the RedPajama Dataset. On the RedPajama Dataset we find, at worst, a 0.3 bits per byte increase in KL divergence to the models final output. On the helpful harmless corpus we find no significant difference between the transferred and trained lenses. These results show that fine-tuning minimally effects the representations used by the tuned lens. This opens applications of the tuned lens for monitoring changes in the representations of a module during fine-tuning and minimizes the need for practitioners to train lenses on new fine-tuned models.
@@ 4.1 Causal basis extraction
[…]
@@ 4.2 Stimulus-response alignment
[…]
@@
Given a hidden state $\boldsymbol{h}_{\ell}=\mathcal{M}_{\leq\ell}(\boldsymbol{x})$, resampling ablation replaces the principal subspace of $\boldsymbol{h}_{\ell}$ with the corresponding subspace generated on a different input $\boldsymbol{x}^{\prime}$ selected uniformly at random from the dataset. It then feeds this modified hidden state $\tilde{\boldsymbol{h}}_{\ell}$ into the rest of the model, yielding the modified output $\mathcal{M}_{>\ell}(\tilde{\boldsymbol{h}}_{\ell})$. Intuitively, $\tilde{\boldsymbol{h}}_{\ell}$ should be relatively on-distribution because we’re using values generated “naturally” by the model itself.
@@ 5.2 Detecting Prompt Injections
ARC-Easy $0.59\;(0.54,0.62)$ $0.73\;(0.71,0.76)$ $0.53\;(0.50,0.57)$ $0.59\;(0.56,0.62)$ $0.73\;(0.70,0.75)$ $72.8\%\rightarrow 31.7\%$
ARC-Challenge $0.71\;(0.65,0.77)$ $0.81\;(0.77,0.84)$ $0.73\;(0.67,0.79)$ $0.80\;(0.77,0.83)$ $0.57\;(0.53,0.61)$ $43.5\%\rightarrow 24.7\%$
BoolQ $0.99\;(0.98,0.99)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.89\;(0.87,0.91)$ $0.61\;(0.57,0.66)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $67.1\%\rightarrow 0.0\%$
MC TACO $0.74\;(0.71,0.77)$ $0.68\;(0.66,0.70)$ $0.68\;(0.66,0.69)$ $0.55\;(0.53,0.59)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.40\rightarrow 0.06$ F1
MNLI $0.98\;(0.98,0.99)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.95\;(0.94,0.96)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $54.3\%\rightarrow 0.0\%$
QNLI $0.99\;(0.99,1.00)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.93\;(0.92,0.95)$ $0.68\;(0.63,0.71)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $54.3\%\rightarrow 0.0\%$
QQP $1.00\;(0.99,1.00)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.90\;(0.89,0.90)$ $0.79\;(0.76,0.81)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $60.7\%\rightarrow 6.5\%$
SciQ $0.62\;(0.57,0.69)$ $0.64\;(0.59,0.70)$ $0.75\;(0.71,0.79)$ $0.70\;(0.65,0.74)$ $0.75\;(0.72,0.78)$ $95.5\%\rightarrow 62.6\%$
SST-2 $1.00\;(0.98,1.00)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $0.78\;(0.72,0.83)$ $0.61\;(0.56,0.65)$ 1.00(1.00,1.00)$1.00$\;(1.00,1.00)1.00 ( 1.00 , 1.00 ) $82.9\%\rightarrow 49.1\%$
@@ 5.3 Measuring Example Difficulty
QNLI $\mathbf{0.409}$ $-0.099$ $53.0\%$
QQP $\mathbf{0.585}$ $-0.340$ $0.381$ (F1)
@@ 6 Discussion
Limitations and future work. One limitation of our method is that it involves training a translator layer for each layer of the network, while the logit lens can be used on any pretrained model out-of-the-box. This training process, however, is quite fast: our code can train a full set of probes in under an hour on a single 8$\times$×A40 node, and further speedups are likely possible. We have also released tuned lens checkpoints for the most commonly used pretrained models as part of our tuned-lens library, which should eliminate this problem for most applications.
@@ Appendix C Transformers Perform Iterative Inference
We reproduced their experiment using Pythia 6.9B, and report the results in Figure 19. We show that, for every layer in the network, the cosine similarity between the residual and the gradient is negative at least 95% of the time. While the magnitudes of the cosine similarities are relatively small in absolute terms, never exceeding 0.05, we show that they are much larger than would be expected of random vectors in this very high dimensional space. Specifically, we first sample 250 random Gaussian vectors of the same dimensionality as $\frac{\partial\mathcal{L}(\boldsymbol{h}_{i})}{\partial\boldsymbol{h}_{i}}$, which is (hidden size) $\times$ (sequence length) $=8{,}388{,}608$. We then compute pairwise cosine similarities between the vectors, and find the 5th percentile of this sample to be $\mathbf{-6\times 10^{-4}}$. Virtually all of the gradient-residual pairs we observed had cosine similarities below this number.
@@ E.1 Static Analysis
Table 3: The mean interpretability scores (as measured in Section E.3) for Pythia 125M, with several different interpretability techniques [Millidge and Black, 2022, Geva et al., 2020], comparing both the tuned lens and logit lens to randomly generated matrices. Where applicable, the notation (L) and (R) indicates that the results are for the left and right singular vectors, respectively.
@@ E.5 Implementation
VN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v5

@@ 1 Introduction
We use the tuned lens to gain qualitative insight into the computational process of pretrained language models, by examining how their latent predictions evolve during a forward pass (Figure˜1, Appendix˜B)
@@ 2 The Logit Lens
$\mathcal{M}_{>\ell}(\boldsymbol{h}_{\ell})=\mathrm{LayerNorm}\Big{[}\boldsymbol{h}_{\ell}+\sum_{\ell^{\prime}=\ell}^{L}\underbrace{F_{\ell^{\prime}}(\boldsymbol{h}_{\ell^{\prime}})}_{\text{residual update}}\hskip-5.78172pt\Big{]}W_{U}.$ (2)
@@
Moreover, the type of information extracted by the logit lens varies both from model to model and from layer to layer, making it difficult to interpret. For example, we find that for BLOOM and OPT 125M, the top 1 prediction of the logit lens is often the input token, rather than any plausible continuation token, in more than half the layers (Figure˜18).
@@
$\forall v\in\mathcal{V},\quad\mathcal{V}$ $=\{\texttt{"aardvark"},\ldots\}$
@@ 3 The Tuned Lens
Figure 6: Pairwise similarities of hidden state covariance matrices across layers of Pythia 12B. Layer 4 introduces two outlier dimensions which dominate the covariance; removing them reveals smooth representational drift with depth. To control for varying hidden state norms, we measure the Frobenius cosine similarity, or $\frac{\langle A,B\rangle_{F}}{\|A\|_{F}\|B\|_{F}}$ for two matrices AA and $B$.
Results. We plot tuned lens perplexity as a function of depth for the Pythia models and GPT-NeoX-20B in Figure˜5333Pythia and GPT-NeoX-20B were trained using the same architecture, data, and codebase (Andonian et al., 2021). While they’re not officially the same model suite, they’re more consistent than the OPT models.; results for other model families can be found in Appendix A.
We find that the tuned lens resolves the problems with the logit lens discussed in Section 2: it has significantly lower bias (Figure˜3), and much lower perplexity than the logit lens across the board (Figure˜5, Appendix A).
@@
We report transfer penalties for the largest Pythia model in Figure˜7. Overall, transfer penalties are quite low, especially for nearby layers (entries near the diagonal in Figure˜7). Comparing to the two plots in Figure˜6, we notice that transfer penalties are strongly negatively correlated with covariance similarity (Spearman $\rho=-0.78$). Unlike Figure˜6, however, Figure˜7 is not symmetric: transfer penalties are higher when training on a layer with the outlier dimensions (Layer 5 and later) and testing on a layer without them, than the reverse.
@@
As an example of the transferability of lenses to fine-tuned models we use Vicuna 13B (Chiang et al., 2023), an open source instruct fine-tuned chat model based on LLaMA 13B (Touvron et al., 2023). In Figure˜12, Appendix˜A, we compare the performance of a lens specifically trained on Vicuna to a tuned lens trained only on LLaMA. Both Tuned Lenses were trained using a subsample of the RedPajama dataset (Together, 2023), an open source replication of LLaMA’s training set. The performance of both lenses was then evaluated on the test set of Anthropic’s Helpful Harmless (Bai et al., 2022) conversation dataset and the RedPajama Dataset. On the RedPajama Dataset we find, at worst, a 0.3 bits per byte increase in KL divergence to the models final output. On the helpful harmless corpus we find no significant difference between the transferred and trained lenses. These results show that fine-tuning minimally effects the representations used by the tuned lens. This opens applications of the tuned lens for monitoring changes in the representations of a module during fine-tuning and minimizes the need for practitioners to train lenses on new fine-tuned models.
@@ 4.1 Causal basis extraction
$\textrm{s.t.}\quad\langle\boldsymbol{v},\boldsymbol{v}_{j}\rangle$ $=\boldsymbol{0},\quad\forall j<i$
@@ 4.2 Stimulus-response alignment
$\langle\mathbf{p}_{\mathrm{new}}-\mathbf{p}_{\mathrm{old}},\;\mathbf{q}_{\mathrm{new}}-\mathbf{q}_{\mathrm{old}}\rangle_{\mathbf{w}}\;>\;0.$ (15)
@@
[…]
@@ 5.2 Detecting Prompt Injections
ARC-Easy 0.59(0.54,0.62)0.59\;(0.54,0.62) 0.73(0.71,0.76)0.73\;(0.71,0.76) 0.53(0.50,0.57)0.53\;(0.50,0.57) 0.59(0.56,0.62)0.59\;(0.56,0.62) 0.73(0.70,0.75)0.73\;(0.70,0.75) $72.8\%\rightarrow 31.7\%$
ARC-Challenge 0.71(0.65,0.77)0.71\;(0.65,0.77) 0.81(0.77,0.84)0.81\;(0.77,0.84) 0.73(0.67,0.79)0.73\;(0.67,0.79) 0.80(0.77,0.83)0.80\;(0.77,0.83) 0.57(0.53,0.61)0.57\;(0.53,0.61) $43.5\%\rightarrow 24.7\%$
BoolQ 0.99(0.98,0.99)0.99\;(0.98,0.99) 1.00(1.00,1.00)1.00\;(1.00,1.00) 0.89(0.87,0.91)0.89\;(0.87,0.91) 0.61(0.57,0.66)0.61\;(0.57,0.66) 1.00(1.00,1.00)1.00\;(1.00,1.00) $67.1\%\rightarrow 0.0\%$
MC TACO 0.74(0.71,0.77)0.74\;(0.71,0.77) 0.68(0.66,0.70)0.68\;(0.66,0.70) 0.68(0.66,0.69)0.68\;(0.66,0.69) 0.55(0.53,0.59)0.55\;(0.53,0.59) 1.00(1.00,1.00)1.00\;(1.00,1.00) $0.40\rightarrow 0.06$ F1
MNLI 0.98(0.98,0.99)0.98\;(0.98,0.99) 1.00(1.00,1.00)1.00\;(1.00,1.00) 0.95(0.94,0.96)0.95\;(0.94,0.96) 1.00(1.00,1.00)1.00\;(1.00,1.00) 1.00(1.00,1.00)1.00\;(1.00,1.00) $54.3\%\rightarrow 0.0\%$
QNLI 0.99(0.99,1.00)0.99\;(0.99,1.00) 1.00(1.00,1.00)1.00\;(1.00,1.00) 0.93(0.92,0.95)0.93\;(0.92,0.95) 0.68(0.63,0.71)0.68\;(0.63,0.71) 1.00(1.00,1.00)1.00\;(1.00,1.00) $54.3\%\rightarrow 0.0\%$
QQP 1.00(0.99,1.00)1.00\;(0.99,1.00) 1.00(1.00,1.00)1.00\;(1.00,1.00) 0.90(0.89,0.90)0.90\;(0.89,0.90) 0.79(0.76,0.81)0.79\;(0.76,0.81) 1.00(1.00,1.00)1.00\;(1.00,1.00) $60.7\%\rightarrow 6.5\%$
SciQ 0.62(0.57,0.69)0.62\;(0.57,0.69) 0.64(0.59,0.70)0.64\;(0.59,0.70) 0.75(0.71,0.79)0.75\;(0.71,0.79) 0.70(0.65,0.74)0.70\;(0.65,0.74) 0.75(0.72,0.78)0.75\;(0.72,0.78) $95.5\%\rightarrow 62.6\%$
SST-2 1.00(0.98,1.00)1.00\;(0.98,1.00) 1.00(1.00,1.00)1.00\;(1.00,1.00) 0.78(0.72,0.83)0.78\;(0.72,0.83) 0.61(0.56,0.65)0.61\;(0.56,0.65) 1.00(1.00,1.00)1.00\;(1.00,1.00) $82.9\%\rightarrow 49.1\%$
@@ 5.3 Measuring Example Difficulty
QNLI $\mathbf{0.409}$ $0.099-$ $53.0\%$
QQP $\mathbf{0.585}$ $0.340-$ $0.381$ (F1)
@@ 6 Discussion
[…]
@@ Appendix C Transformers Perform Iterative Inference
We reproduced their experiment using Pythia 6.9B, and report the results in Figure 19. We show that, for every layer in the network, the cosine similarity between the residual and the gradient is negative at least 95% of the time. While the magnitudes of the cosine similarities are relatively small in absolute terms, never exceeding 0.05, we show that they are much larger than would be expected of random vectors in this very high dimensional space. Specifically, we first sample 250 random Gaussian vectors of the same dimensionality as $\frac{\partial\mathcal{L}(\boldsymbol{h}_{i})}{\partial\boldsymbol{h}_{i}}$, which is (hidden size) $\times$ (sequence length) =8,388,608=8{,}388{,}608. We then compute pairwise cosine similarities between the vectors, and find the 5th percentile of this sample to be $\mathbf{-6\times 10^{-4}}$. Virtually all of the gradient-residual pairs we observed had cosine similarities below this number.
@@ E.1 Static Analysis
Table 3: The mean interpretability scores (as measured in Section˜E.3) for Pythia 125M, with several different interpretability techniques [Millidge and Black, 2022, Geva et al., 2020], comparing both the tuned lens and logit lens to randomly generated matrices. Where applicable, the notation (L) and (R) indicates that the results are for the left and right singular vectors, respectively.